OpenStudy (anonymous):

given f(x) = x^2 + 4. for x and y are real positif number thas satisfying f(xy) + f(y − x) = f(y + x). find the maximum value of x + y

OpenStudy (anonymous):

@rational

OpenStudy (rational):

IMO problem again ha

OpenStudy (anonymous):

no it is not :D

OpenStudy (anonymous):

my work is : f(xy) = f(y + x) - f(y -x) (xy)^2 + 4 = (y+x)^2 + 4 -((y-x)^2 + 4) (xy)^2 + 4 = 2xy + 2xy (xy)^2 + 4 = 4xy (xy)^2 - 4xy + 4 = 0 (xy - 2)^2 = 0 xy = 2

OpenStudy (anonymous):

what next ?

OpenStudy (anonymous):

set x = y ?

OpenStudy (rational):

looks neat !!!!

OpenStudy (rational):

yes that will do

OpenStudy (rational):

maximize x+y subject to xy=2

OpenStudy (anonymous):

so, max of (x+y) = 2 sqrt(2) ?

OpenStudy (rational):

yes that should work :)

OpenStudy (anonymous):

xy=2 ---> y = 2/x k = x + y k = x + 2/x k = (x^2 + 2)/x k' = (2x(x) - 1(x^2 + 2))/x^2 = 0 2x^2 - x^2 - 2 = 0 x = (1 +- sqrt((-1)^2 - 4(2)(-2))/(2(2)) x = (1 + sqrt(17)/4 k max = ((1 + sqrt(17)/4)^2 + 2)/ ((1 + sqrt(17)/4

OpenStudy (anonymous):

hmm... i think there is something wrong with my job

OpenStudy (rational):

``` xy=2 ---> y = 2/x k = x + y k = x + 2/x k = (x^2 + 2)/x k' = (2x(x) - 1(x^2 + 2))/x^2 = 0 2x^2 - x^2 - 2 = 0 ``` immediately after this, dont you get `x^2 - 2 = 0` ?

OpenStudy (anonymous):

opsss.. yes that's my mistake. that's should be : x^2 - 2 = 0 x^2 = 2 x = +- sqrt(2) obvious take x positive, so k max = (x^2 + 2)/x = (2 + 2)/sqrt(2) = 2sqrt(2) thanks :D

OpenStudy (anonymous):

but i dont understand why the shortway is just take x = y ?

OpenStudy (rational):

thats a geometric argument, given that area of a rectangle = xy = 2 the largest possible perimeter is achieved for a square shape

OpenStudy (anonymous):

oh, ok.. thank you very much :D

OpenStudy (rational):

np :)

OpenStudy (anonymous):

wait, why the wolfram said max is infinite ? --"

OpenStudy (anonymous):

is the question should asking us find the minimum value ?

OpenStudy (rational):

Ahh yes it seems the problem is more involved

OpenStudy (freckles):

you probably know this but i think you made that differentiating task harder than it should have been

OpenStudy (freckles):

\[k=x+\frac{2}{x} \\ k'=1-\frac{2}{x^2} \\ 1-\frac{2}{x^2}=0 \\ 1=\frac{2}{x^2} \\ x^2=2\]